OpenStudy (anonymous):

If the Fahrenheit measure of temperature is a function of the Celsius measure, and the Fahrenheit measure is 32 more than 9/5 times the Celsius measure how do I use g to represent the function?

OpenStudy (anonymous):

Is this just asking for the Fahrenheit formula in terms of C?

OpenStudy (anonymous):

I'm sorry, I really don't know

OpenStudy (anonymous):

I'm not 100% sure, but that what it seems to me. Wait for someone to confirm. Do you know How to solve for F in term of C?

OpenStudy (anonymous):

I believe he wants us to use variables like independent, dependent etc. I just don't understand that

OpenStudy (anonymous):

and no to your answer

OpenStudy (anonymous):

Do you not understand dependent an independent all together or just for this specific example

OpenStudy (anonymous):

altogether

OpenStudy (skullpatrol):

You are right, the question asks for you to assign letters to the variables :-)

OpenStudy (skullpatrol):

Let F = the Fahrenheit measure of temperature Let C = the Celsius measure of temperature

OpenStudy (anonymous):

If we graph distance as function of time, this means that on the x axis we would put time and on the y axis we would put distance. So let's say after 1 hour we drove 60 miles, at point x = 1, y = 60 at point x = 2 y = 120. X which is time is independent. The distance traveled is dependent on time. Does this make sense?

OpenStudy (anonymous):

Another way of putting it; time is not dependent on distance, rather distance is dependent on time.

OpenStudy (anonymous):

yes that makes sense

OpenStudy (skullpatrol):

Can you translate the the words in the sentence into an equation?

OpenStudy (anonymous):

no skullpatrol

OpenStudy (anonymous):

So the x axis is independent and y axis is dependent. Another way of thinking about it is x = is input y = output. If you can graph out the following on paper it would be helpful...

OpenStudy (skullpatrol):

F = 32 + (9/5)*C

OpenStudy (anonymous):

Do you know the freezing and boiling points of water in F and C? I'm trying to help you understand why the formula that was just posted is true.

OpenStudy (anonymous):

skullpatrol I do get that or it could be1.8 instead of 9/5 right?

OpenStudy (skullpatrol):

To use function notation: F = g(C) = (9/5)*C + 32

OpenStudy (skullpatrol):

Yes, 9/5 = 1.8

OpenStudy (skullpatrol):

Answer: g(C) = 1.8*C + 32

OpenStudy (anonymous):

THANK YOU SO MUCH!!! i GET IT!!

OpenStudy (skullpatrol):

Note: F = g(C) is a symbolic way of saying the temperature as measured in Fahrenheit is a function, g of the temperature measured in Celsius.

OpenStudy (anonymous):

LEIBY i BELIEVE IT IS 0 C and 32F for freezing and 212F and 100C for boiling

OpenStudy (anonymous):

Correct. If you want do understand why F = 32 + (9/5)*C. Let me know and I'll show you.

OpenStudy (anonymous):

Please show me

OpenStudy (anonymous):

Since water freezes at 0 degrees C and boils at 100 degrees C. Draw on the x axis a line from 0 to 100. (It doesn't need to be exact, you can make markings at increments of 10 if you'd like.)

OpenStudy (anonymous):

Make the Y axis go up till 220. Mark on the Y axis where the freezing point and boiling points are (32 and 212). Make a line between the 2 points. Since y = slope x + intercept, calculate the slope which is (y2 - y1) / (x2 - x1) which is (212 - 32)/ 100 - 0) which is 180 / 100 simplified = 9/5 so the slpe is 9/5. The intercept is 32. So, Y = (9/5)x + 32. Or F = (9/5)C + 32

OpenStudy (anonymous):

wow!! you are brilliant! Math is not my subject but I keep trying, now that I found open study I may get pretty good at it with people like you and skull. Thank you sooooo much for your patience and understanding. Now I am going to write down what you said and go to bed!!! I am sooo tired.

OpenStudy (anonymous):

I'm not brilliant, I also keep trying, till I get it.

OpenStudy (anonymous):

you have a nice rest of the night! Well, it is after 1 am now! Gotta get up at 6:30. What state do you live in?

OpenStudy (anonymous):

One more thing, so which would be the independent and the dependent variable?

OpenStudy (anonymous):

x is independent y is dependent So in F = 32 + (9/5)*C F is dependent on C

OpenStudy (anonymous):

got it!! Thank you again!